Find the gain and phase margin when the plant and controller are: P(s) = 1/(s+1), C(s) = 1/s
Answer #1

G(s) = P(s). Cos) (5+1) S S(5+1) O phase margin 180+ phase Angle (8) = 2 Lulwa W = Wgc Gain crops over trequency (wach lacjw)

180 + phare margin = 180- 121.72 : 58.38 Gain margin (Gm) | aju) M(jw) Gm= 1/4 I waupe wpc (phase cross over frequency) Lucj